# Recognition of Unit Segment and Polyline Graphs is $\exists\mathbb{R}$-Complete

### v1

@@ 1 Introduction
Many real-life problems can be mathematically described in the language of graphs. For instance, Cellnex Telecom owns more than $2000$ cell towers in Switzerland. We want to assign each tower a frequency such that no two towers that overlap in coverage use the same frequency. This becomes a graph coloring problem. Every cell tower becomes a vertex, overlap indicates an edge and a frequency assignment corresponds to a proper coloring of the vertices, see Figure 1.
@@
In many contexts, we have additional structure on the graph that may or may not help us to solve the underlying algorithmic problem. For instance, it might be that the graph arises as the intersection graph of unit disks in the plane (each unit disk gives a vertex, and two vertices are adjacent if their corresponding disks overlap). In that case, the coloring problem can be solved more efficiently [10], and there are better approximation algorithms for the clique problem [11]. This motivates a systematic study of geometric intersection graphs. Another motivation is mathematical curiosity. Simple geometric shapes are easily visualized and arguably very natural mathematical objects. Studying the structural properties of intersection graphs gives insight into those geometric shapes and their possible intersection patterns.
It is known for a host of geometric shapes that it is $\exists\mathbb{R}$-complete to recognize their intersection graphs [28, 14, 25, 27]. The class $\exists\mathbb{R}$ consists of all of those problems that are polynomial-time equivalent to deciding whether a polynomial $p\in\mathbb{Z}[X_{1},\ldots,X_{n}]$ has a root. We will introduce $\exists\mathbb{R}$ in more detail below. $\exists\mathbb{R}$-completeness is known for the recognition problems of intersection graphs for segments, disks, unit disks, rays, grounded segments, downward rays, and a few other examples.
@@ Results.
PolyLine Recognition is $\exists\mathbb{R}$-complete, for any fixed $k\geq 1$.
@@ 1.2 Discussion
We see our main contribution in refining the hierarchy of geometric graph classes for which recognition complexity is known. Both unit segments as well as polylines with $k$ bends are natural objects that are well studied in the literature. However, the recognition of the corresponding graph classes was not studied previously. Polylines with an unbounded number of bends are equivalent to strings [^1] It is possible to show that polylines with an unbounded number of bends are as versatile as strings with respect to the types of graphs that they can represent, since the number of intersections of any two strings can always be bounded from above [36, 37]., while polylines with 00 bends are just segments. Polylines with $k$ bends thus naturally slot in between strings and segments, and their corresponding graph class is thus also an intermediate class between the class of segment intersection graphs and string graphs, as can be seen in Figure 2. By showing that recognition for polylines with $k$ bends is $\exists\mathbb{R}$-complete for all constant $k$, we see that the switch from $\exists\mathbb{R}$-completeness (segment intersection graphs) to NP-membership (string graphs) really only happens once $k$ is infinite. Similarly, unit segment intersection graphs slot in between segment and ray intersection graphs. Intuitively, recognition of a class intermediate to two classes that are $\exists\mathbb{R}$-hard to recognize should also be $\exists\mathbb{R}$-hard, and our Theorem 1 confirms this intuition in this case.
@@
One of the consequences of $\exists\mathbb{R}$-completeness is that there are no short representations of solutions known. Some representable graphs may only be representable by objects with irrational coordinates, or by rational coordinates with nominator and denominator of size at least $2^{2^{n^{c}}}$, for some fixed $c>0$. In other words, the numbers to describe the position might need to be doubly exponentially large [27] for some graphs. For “flexible” objects like polylines, rational solutions can always be obtained by slightly perturbing the representation, however for more “sturdy” objects like unit segments this may not be possible.
@@
The class of the existential theory of the reals $\exists\mathbb{R}$ (pronounced as ‘ER’) is a complexity class which has gained a lot of interest, specifically within the computational geometry community. To define this class, we first consider the problem ETR, which also stands for Existential Theory of the Reals. In an instance of this problem we are given a sentence of the form
$\exists x_{1},\ldots,x_{n}\in\mathbb{R}:\Phi(x_{1},\ldots,x_{n}),$
where $\Phi$ is a well-formed quantifier-free formula consisting of the symbols $\{0,1,x_{1},\ldots,x_{n},+,\cdot,\geq,>,\wedge,\vee,\neg\}$, the goal is to check whether this sentence is true. As an example of an ETR-instance, we could take $\Phi=x\cdot y^{2}+x\geq 0\wedge\neg(y<2x)$. The goal of this instance would be to determine whether there exist real numbers $x$ and $y$ satisfying this formula.
Now the class $\exists\mathbb{R}$ is the family of all problems that admit a polynomial-time many-one reduction to ETR. It is known that
$\textsf{NP}\subseteq\exists\mathbb{R}\subseteq\textsf{PSPACE}.$
The first inclusion follows from the definition of $\exists\mathbb{R}$ as follows. Given any Boolean satisfiability formula, we can replace each positive occurrence of a variable $x$ by $x=1$. For example $(x\lor\lnot y)\land(\lnot x\lor z)$ becomes $(x=1\lor\lnot y=1)\land(\lnot x=1\lor z=1)$.
Showing the second inclusion was first done by Canny in his seminal paper [13]. The reason that $\exists\mathbb{R}$ is an important complexity class is that a number of common problems, mainly in computational geometry, have been shown to be complete for this class, see below.
We want to point out that there are some subtleties in the definition of ETR. Since a sentence in ETR can only contain the integers 00 and $1$, it naively takes $\Theta(n)$ bits to encode a natural number $n$, i.e., $n=1+1+\ldots+1$. However, it is possible to encode $n$ in $O(\log n)$ bits using Horner’s Schema (i.e., $15=1+2\cdot 7=1+2(1+6)=\ldots$). This makes ETR as defined above just as powerful as a definition that would also allow for arbitrary integers encoded in binary. Furthermore, we want to emphasize that the reductions used for defining $\exists\mathbb{R}$ are performed in the word RAM model (or equivalently on a Turing machine), and not on a real Random Access Machine (real RAM) or in the Blum-Shub-Smale model.
The main reason that the complexity class $\exists\mathbb{R}$ gained traction in recent years is the increasing number of important algorithmic problems that are $\exists\mathbb{R}$-complete. Schaefer established the current name and pointed out first that several known NP-hardness reductions actually imply $\exists\mathbb{R}$-completeness [33]. Note that some important reductions that establish $\exists\mathbb{R}$-completeness were done before the class was named.
Problems that have a continuous solution space and non-linear relation between partial solutions are natural candidates to be $\exists\mathbb{R}$-complete. Early examples are related to recognition of geometric structures: points in the plane [29, 42], geometric linkages [34, 1], segment graphs [25, 27], unit disk graphs [28, 21], ray intersection graphs [14], and point visibility graphs [14]. In general, the complexity class is more established in the graph drawing community [26, 16, 35, 18]. Yet, it is also relevant for studying polytopes [31, 17]. There is a series of papers related to Nash-Equilibria [6, 38, 20, 8, 9]. Another line of research studies matrix factorization problems [15, 40, 41, 39]. Other $\exists\mathbb{R}$-complete problems are the Art Gallery Problem [3, 44], covering polygons with convex polygons [2], geometric packing [5] and training neural networks [4, 7].
@@ 1.4 Proof Techniques
Figure 3: The pseudoline arrangement on the left is combinatorially equivalent to the line arrangement on the right and thus stretchable.
The techniques used in this paper are similar to previous work. In essence, we are reducing from the Stretchability problem. In this problem, we are given a pseudoline arrangement as an input, and the question is whether this arrangement is stretchable (see below for a detailed explanation). Given the initial pseudoline arrangement, we construct a graph. Certain vertices represent the pseudolines, and in any representation of the graph by unit segments or polylines their realizations witness that the pseudolines are stretchable. We use a long cycle surrounding the construction to enforce a certain order on any representation. Furthermore, we use probes to guarantee the desired combinatorial structure. Both general ideas have already been used in different contexts [14].
The main idea to show hardness of PolyLine Recognition is to build a construction where the polylines have to make $k$ bends such that the crucial part has no bends and witnesses the stretchability of the original pseudoline arrangement.
Definition 3 (Pseudoline arrangement).
A pseudoline arrangement $\mathcal{A}$ is a set of $n$ curves that are $x$-monotone. Furthermore, any two curves intersect exactly once and no three curves meet in a single point. We assume that there exist two vertical lines on which each curve starts and ends. No intersections of curves lie on these vertical lines.
Note that the literature sometimes distinguishes between simple and general pseudoline arrangements. For us all the pseudoline arrangements are simple, so we drop the term simple, similarly to how we speak of graphs meaning simple graphs. Further note that line arrangements that are truncated far enough are also pseudoline arrangements.
We say two pseudoline arrangements have the same combinatorics if there exists an isomorphism (bijective and continuous) of the plane that deforms one into the other. It is known that the order in which each pseudoline intersects the remaining pseudolines determines the combinatorial type completely. We say a pseudoline arrangement is stretchable, if there exists a line arrangement with the same combinatorics. Compare to Figure 3.
@@ 2.2 Order Lemma
3 Proofs
In this section, we first show the $\exists\mathbb{R}$-membership parts of Theorems 1 and 2. Then, we prove $\exists\mathbb{R}$-hardness, first for unit segments, then for polylines. The reduction for polylines builds upon the reduction for unit segments and we will only highlight the differences.
3.1 $\exists\mathbb{R}$-Membership
@@
Given this alternative characterization of $\exists\mathbb{R}$, it is now very easy to establish $\exists\mathbb{R}$-membership of Unit Recognition and PolyLine Recognition: We merely need to describe the witness and the verification algorithm. The witness is a description of the coordinates of the unit segments (or polylines, respectively) realizing the given graph. The verification algorithm merely checks that the correct realizations of vertices intersect one another, and in the case of unit segments, also checks that all segments have the correct length. Note that we sweep many details of the algorithm under the carpet. However, algorithms are much more versatile than formulas and it is a well-established fact that algorithms are capable of all types of elementary operations needed to perform this verification.
@@ 3.2 $\exists\mathbb{R}$-Hardness for Unit Segments
This section is dedicated to show $\exists\mathbb{R}$-hardness of Unit Recognition by a reduction from Stretchability.
@@
Given a pseudoline arrangement $\mathcal{A}$ of $n$ pseudolines as defined in Definition 3, we construct a graph $G$ by enhancing the arrangement $\mathcal{A}$ with additional Jordan arcs. Then we define $G$ to be the intersection graph of the pseudolines (which are also Jordan arcs) and all our added arcs. See Figure 8 for an illustration.
@@
Figure 11: We can place all the connector segments correctly on the left and on the right using a simple sawtooth pattern.
@@ 3.3 $\exists\mathbb{R}$-Hardness for Polylines
In this section, we study polylines with $k$ bends and we assume $k>0$ is a fixed constant. Specifically, we will show that PolyLine Recognition is $\exists\mathbb{R}$-complete.

### v2

@@ 1 Introduction
Many real-life problems can be mathematically described in the language of graphs. For instance, consider all the cell towers in Switzerland. We want to assign each tower a frequency such that no two towers that overlap in coverage use the same frequency. This can be seen as a graph coloring problem: Every cell tower becomes a vertex, overlap indicates an edge and a frequency assignment corresponds to a proper coloring of the vertices, see Figure 1.
@@
In many contexts, we have additional structure on the graph that may or may not help us to solve the underlying algorithmic problem. For instance, it might be that the graph arises as the intersection graph of unit disks in the plane such as in the example above. In that case, the coloring problem can be solved more efficiently [10], and there are better approximation algorithms for the clique problem [11]. This motivates a systematic study of geometric intersection graphs. Another motivation is mathematical curiosity. Simple geometric shapes are easily visualized and arguably very natural mathematical objects. Studying the structural properties of intersection graphs gives insight into those geometric shapes and their possible intersection patterns.
It is known for a host of geometric shapes that it is $\exists\mathbb{R}$-complete to recognize their intersection graphs [14, 25, 27, 28]. The class $\exists\mathbb{R}$ consists of all of those problems that are polynomial-time equivalent to deciding whether a given system of polynomial inequalities and equations with integer coefficients has a real solution. We will introduce $\exists\mathbb{R}$ in more detail below.
@@ Results.
$k$-PolyLine Recognition is $\exists\mathbb{R}$-complete, for any fixed $k\geq 1$.
@@ 1.2 Discussion
We see our main contribution in refining the hierarchy of geometric graph classes for which recognition complexity is known. Both unit segments as well as polylines with $k$ bends are natural objects that are well studied in the literature. However, the recognition of the corresponding graph classes was not studied previously. Polylines with an unbounded number of bends are equivalent to strings (it is possible to show that polylines with an unbounded number of bends are as versatile as strings with respect to the types of graphs that they can represent, since the number of intersections of any two strings can always be bounded from above [36, 37]), while polylines with 00 bends are just segments. Polylines with $k$ bends thus naturally slot in between strings and segments, and their corresponding graph class is thus also an intermediate class between the class of segment intersection graphs and string graphs, as can be seen in Figure 2. By showing that recognition for polylines with $k$ bends is $\exists\mathbb{R}$-complete for all constant $k$, we see that the switch from $\exists\mathbb{R}$-completeness (segment intersection graphs) to NP-membership (string graphs) really only happens once $k$ is infinite. Similarly, unit segment intersection graphs slot in between segment and ray intersection graphs. Intuitively, recognition of a class intermediate to two classes that are $\exists\mathbb{R}$-hard to recognize should also be $\exists\mathbb{R}$-hard, and our Theorem 1 confirms this intuition in this case.
@@
One of the consequences of $\exists\mathbb{R}$-completeness is that there are no short representations of solutions known. Some representable graphs may only be representable by objects with irrational coordinates, or by rational coordinates with nominator and denominator of size at least $2^{2^{n^{c}}}$, for some fixed $c>0$. In other words, the numbers to describe the position might need to be doubly exponentially large [27] for some graphs. For “flexible” objects like polylines, rational solutions can always be obtained by slightly perturbing the representation. For more “sturdy” objects like unit segments this may not be possible, however it is known that for example unit disks admit rational solutions despite their inflexibility [28].
@@
The class of the existential theory of the reals $\exists\mathbb{R}$ (pronounced as ‘ER’) is a complexity class defined through its canonical problem ETR, which also stands for Existential Theory of the Reals. In this problem we are given a sentence of the form $\exists x_{1},\ldots,x_{n}\in\mathbb{R}:\Phi(x_{1},\ldots,x_{n})$, where $\Phi$ is a well-formed quantifier-free formula consisting of the symbols $\{0,1,x_{1},\ldots,x_{n},+,\cdot,\geq,>,\wedge,\vee,\neg\}$, and the goal is to check whether this sentence is true.
The class $\exists\mathbb{R}$ is the family of all problems that admit a polynomial-time many-one reduction to ETR. It is known that $\textsf{NP}\subseteq\exists\mathbb{R}\subseteq\textsf{PSPACE}$ [13]. The reason that $\exists\mathbb{R}$ is an important complexity class is that a number of common problems, mainly in computational geometry, have been shown to be complete for this class. Schaefer established the current name and pointed out first that several known NP-hardness reductions actually imply $\exists\mathbb{R}$-completeness [33]. Early examples are related to recognition of geometric structures: points in the plane [29, 42], geometric linkages [1, 34], segment graphs [25, 27], unit disk graphs [21, 28], ray intersection graphs [14], and point visibility graphs [14]. In general, the complexity class is more established in the graph drawing community [16, 18, 26, 35]. Yet, it is also relevant for studying polytopes [17, 31], Nash-Equilibria [6, 8, 9, 20, 38], and matrix factorization problems [15, 39, 40, 41]. Other $\exists\mathbb{R}$-complete problems are the Art Gallery Problem [3, 44], covering polygons with convex polygons [2], geometric packing [5] and training neural networks [4, 7].
@@ 1.4 Proof Techniques
Figure 3: The pseudoline arrangement on the left is combinatorially equivalent to the (truncated) line arrangement on the right; hence, it is stretchable.
The techniques used in this paper are similar to previous work. $\exists\mathbb{R}$-membership can be established straightforwardly by constructing concrete formulae or invoking a characterization of $\exists\mathbb{R}$ using real verification algorithms [19], similar to the characterization of NP. This proof can be found in Section 3.
For $\exists\mathbb{R}$-hardness, we are in essence reducing from the Stretchability problem. In this problem, we are given a simple pseudoline arrangement as an input, and the question is whether this arrangement is stretchable. A pseudoline arrangement $\mathcal{A}$ is a set of $n$ curves that are $x$-monotone. Furthermore, any two curves intersect exactly once and no three curves meet in a single point. We assume that there exist two vertical lines on which each curve starts and ends. The problem is to determine whether there exists a combinatorially equivalent (truncated) line arrangement. See Figure 3 for an example.
For the reduction establishing $\exists\mathbb{R}$-hardness of Unit Recognition, we are given a pseudoline arrangement $\mathcal{A}$, and we construct a graph that is representable by unit segments (or polylines with $k$ bends) iff $\mathcal{A}$ is stretchable. This graph is created by enhancing $\mathcal{A}$ with more curves and taking their intersection graph. By the way these additional curves are constructed, they can always be drawn as unit segments in any line arrangement combinatorially equivalent to $\mathcal{A}$, if this line arrangement is first “squeezed” into a canonical form.
For the other direction of the equivalence we need to show that if the graph is representable by unit segments, then $\mathcal{A}$ is stretchable. To help with this proof we establish helper lemmas for both unit segments and polylines stating that cycles in a graph enforce a certain order of intersections of objects around a closed curve in any realization of that graph. The details of these helper lemmas can be found in Section 2.
Using these additional lemmas we then show that the unit segments corresponding to the vertices originally representing our pseudolines must have the same combinatorial structure as our original arrangement $\mathcal{A}$. To ensure this we add so-called probes during the construction of our graph $G$. The ideas of order-enforcing cycles and probes have already been used in different contexts [14].
The main idea to show hardness of $k$-PolyLine Recognition is to enhance the construction for Unit Recognition by an additional area where the polylines representing our pseudolines have to make $k$ bends. This then ensures that the polylines have no bends in the crucial part representing the pseudoline arrangement $\mathcal{A}$, and the same arguments as for unit segments can be used to show that this witnesses the stretchability of the original pseudoline arrangement.
@@ 2.2 Order Lemma
3 $\exists\mathbb{R}$-Membership
In this section, we show the $\exists\mathbb{R}$-membership parts of Theorems 1 and 2.
@@ 3 $\exists\mathbb{R}$-Membership
Given this alternative characterization of $\exists\mathbb{R}$, it is now very easy to establish $\exists\mathbb{R}$-membership of Unit Recognition and $k$-PolyLine Recognition: We merely need to describe the witness and the verification algorithm. The witness is a description of the coordinates of the unit segments (or polylines, respectively) realizing the given graph. The verification algorithm merely checks that the correct realizations of vertices intersect one another, and in the case of unit segments, also checks that all segments have the correct length. Note that we sweep many details of the algorithm under the carpet. However, algorithms are much more versatile than formulas and it is a well-established fact that algorithms are capable of all types of elementary operations needed to perform this verification.
4 $\exists\mathbb{R}$-Hardness
In this section we prove $\exists\mathbb{R}$-hardness, first for unit segments, then for polylines. The reduction for polylines builds upon the reduction for unit segments and we will only highlight the differences.
@@ 4.1 $\exists\mathbb{R}$-Hardness for Unit Segments
We show $\exists\mathbb{R}$-hardness of Unit Recognition by a reduction from Stretchability.
@@
Given a pseudoline arrangement $\mathcal{A}$ of $n$ pseudolines, we construct a graph $G$ by enhancing the arrangement $\mathcal{A}$ with additional Jordan arcs. Then we define $G$ to be the intersection graph of the pseudolines (which are also Jordan arcs) and all our added arcs. See Figure 8 for an illustration.
@@
Figure 11: We can place all the cycle segments correctly on the left and on the right using a simple sawtooth pattern.
@@ 4.2 $\exists\mathbb{R}$-Hardness for Polylines
We study polylines with $k$ bends and we assume $k>0$ is a fixed constant. Specifically, we will show that $k$-PolyLine Recognition is $\exists\mathbb{R}$-complete.